OpenStudy (anonymous):

In a classroom, 5 out 13 students surveyed said they had two or more siblings. If the entire school has 520 students, predict the number of students that have two or more siblings. Write a proportion to model the problem. In a classroom, 5 out 13 students surveyed said they had two or more siblings. If the entire school has 520 students, predict the number of students that have two or more siblings. Write a proportion to model the problem. @Mathematics

OpenStudy (anonymous):

set up proportion: 5/13=x/520 cross multiply: 5*520=13x isolate x: (5*520)/13=x solve: x=200 students